# Results of Fractional Rough Burgers equation in $H^{s}$ space and its application

### v1

@@ 3.2 Well-posedness
Note that $v$ satisfied (3.6), where $f=\frac{1}{2}\partial_{x}(v+X)^{2}$. By Lemma 3.3 and integral by parts, we have
$\|v\|_{L^{2}}^{2}+\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq\|u_{0}\|_{L^{2}}^{2}+2\int_{0}^{t}\|u\|_{L^{2}}^{2}\mathrm{d}t^{\prime}+\int_{0}^{t}\langle v,2\partial_{x}(vX)+\partial_{x}(X^{2})\rangle_{0}\mathrm{d}t^{\prime}.$
By Hölder inequality and Lemma 2.1 and Lemma 2.7, there exist $\alpha\in(\frac{1}{4},\frac{1}{2})$ such that $\frac{\gamma}{2}-\frac{1}{2}>\alpha>1-\frac{\gamma}{2}$ satisfied
$\langle v,2\partial_{x}(vX)+\partial_{x}(X^{2})\rangle_{0}\leq$ $2\|v\|_{H^{\frac{\gamma}{2}}}\|vX\|_{H^{1-\frac{\gamma}{2}}}+\|v\|_{H^{\frac{\gamma}{2}}}\|X^{2}\|_{H^{1-\frac{\gamma}{2}}}$
$\lesssim$ $2\|v\|_{H^{\frac{\gamma}{2}}}^{\frac{2}{\gamma}}\|v\|_{L^{2}}^{2-\frac{2}{\gamma}}\|X\|_{\mathcal{C}^{\alpha}}+\|v\|_{H^{\frac{\gamma}{2}}}\|X\|_{W^{\alpha}}^{2},$
where we use the interpolation $\|f\|_{H^{1-\frac{\gamma}{2}}}\leq C\|f\|_{L^{2}}^{\frac{2\gamma-2}{\gamma}}\|f\|_{H^{\frac{\gamma}{2}}}^{\frac{2-\gamma}{\gamma}}$. Using Young’s inequality, we have
$\|v\|_{L^{2}}^{2}+\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq\|u_{0}\|_{L^{2}}^{2}+\nu\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}+C_{\nu}\int_{0}^{t}(1+\|X\|_{W^{\alpha}}^{\frac{\gamma}{\gamma-1}})\|v\|_{L^{2}}^{2}\mathrm{d}t^{\prime}+\frac{2}{\nu}\|X\|_{C_{T}W^{\alpha}}^{4}t.$ (3.8)
By Gronwall’s inequality, the proposition is proved. ∎
Moreover, next lemma note that we can improve such $v$ into higher regularity space.
@@
Under the assumption in Lemma 3.3, for $v\in C_{T}L^{2}\cap L^{2}_{T}H^{\frac{\gamma}{2}}$, we have energy estimate as follows
$\|\Lambda^{s}v\|_{L^{2}}+2(1-\nu)\int_{0}^{t}\|\Lambda^{s}v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq\|u_{0}\|^{2}_{L^{2}}e^{C_{\nu}\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}}+\int_{0}^{t}e^{C_{\nu}\int_{t^{\prime}}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime\prime}}(\|X\|_{W^{\alpha}}^{2}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+\|X\|_{W^{\alpha}}^{4})\mathrm{d}t^{\prime}$ (3.9)
@@
Recall $\Lambda^{s}v$ satisfied following equation
$\partial_{t}(\Lambda^{s}v)-\Lambda^{\gamma}(\Lambda^{s}v)=\frac{1}{2}\Lambda^{s}\partial_{x}(v^{2})+\Lambda^{s}\partial_{x}(vx)+\frac{1}{2}\Lambda^{s}\partial_{x}(X^{2}).$ (3.10)
Multiply $\Lambda^{s}v$ both sides and integral, we have
$\frac{1}{2}\|\Lambda^{s}v\|_{L^{2}}^{2}+\int_{0}^{t}\|\Lambda^{s}v\|_{\dot{H}^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}=\frac{1}{2}\langle\Lambda^{s}\partial_{x}(v^{2}),\Lambda^{s}v\rangle+\langle\Lambda^{s}\partial_{x}(vX),\Lambda^{s}v\rangle+\frac{1}{2}\langle\Lambda^{s}\partial_{x}(X^{2}),\Lambda^{s}v\rangle.$ (3.11)
For the first term in the right side of (3.11), by Parsavel formula, we have
$\frac{1}{2}\langle\Lambda^{s}\partial_{x}(v^{2}),\Lambda^{s}v\rangle\leq$ $\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}\|D^{{s+1-\frac{\gamma}{2}}}v\|_{L^{2}}$
$\leq$ $\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}\|D^{{s+1-\frac{\gamma}{2}}}v-2vD^{s+1-\frac{\gamma}{2}}\|_{L^{2}}+\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}\|2vD^{s+1-\frac{\gamma}{2}}\|_{L^{2}}.$
By the estimate in [KPV93, GO14] and the interpolation, we have
$\|D^{{s+1-\frac{\gamma}{2}}}v-2vD^{s+1-\frac{\gamma}{2}}\|_{L^{2}}\lesssim$ $\|D^{\alpha}v\|_{L^{4}}^{2}$
@@
for $\gamma\geq\frac{3}{2}$. By Hölder’s inequality and Sobolev embedding in [BCD11], for $0<s+1-\frac{\gamma}{2}<1$ and $\alpha=\frac{s}{2}+\frac{1}{2}-\frac{\gamma}{4}$, we have
$\|2vD^{s+1-\frac{\gamma}{2}}\|_{L^{2}}\lesssim$ $2\|v\|_{L^{\frac{2}{1-2s}}}\|D^{s+1-\frac{\gamma}{2}}v\|_{L^{\frac{2}{2s}}}$
$\lesssim$ $2\|v\|_{\dot{H}^{s}}\|D^{s+1-\frac{\gamma}{2}}v\|_{\dot{H}^{\frac{1}{2}-s}}$
$\lesssim$ $2\|v\|_{\dot{H}^{s}}\|v\|_{\dot{H}^{\frac{3}{2}-\frac{\gamma}{2}}}$
$\lesssim$ $2\|v\|_{\dot{H}^{s}}\|v\|_{\dot{H}^{\frac{\gamma}{2}}},$
@@
$\frac{1}{2}\langle\Lambda^{s}\partial_{x}(v^{2}),\Lambda^{s}v\rangle\leq\frac{\nu}{3}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}^{2}+C_{\nu}\|v\|_{\dot{H}^{s}}^{2}\|v\|_{\dot{H}^{\frac{\gamma}{2}}}^{2}.$
For the second term and third term, by Hölder’s inequality and Lemma 2.1, we have
$\langle\Lambda^{s}\partial_{x}(vX),\Lambda^{s}v\rangle\leq$ $\|vX\|_{\dot{H}^{s+1-\frac{\gamma}{2}}}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}$
$\leq$ $\|X\|_{W^{\alpha}}\|v\|_{\dot{H}^{s+1-\frac{\gamma}{2}}}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}$
$\leq$ $\|X\|_{W^{\alpha}}\|v\|_{\dot{H}^{\frac{\gamma}{2}}}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}$
$\leq$ $C_{\nu}\|X\|_{W^{\alpha}}^{2}\|v\|^{2}_{\dot{H}^{\frac{\gamma}{2}}}+\frac{\nu}{3}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}^{2},$
@@
$\frac{1}{2}\langle\Lambda^{s}\partial_{x}(X^{2}),\Lambda^{s}v\rangle\leq$ $\|vX\|_{\dot{H}^{s+1-\frac{\gamma}{2}}}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}$
@@
Summary above all, we have
$\frac{1}{2}\|\Lambda^{s}v\|_{L^{2}}^{2}+(1-\nu)\int_{0}^{t}\|\Lambda^{s}v\|_{\dot{H}^{\frac{\gamma}{2}}}\leq$ $C_{\nu}\int_{0}^{t}\|X\|_{W^{\alpha}}^{2}\|v\|_{\dot{H}^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}+C_{\nu}\int_{0}^{t}\|X\|_{W^{\alpha}}^{4}\mathrm{d}t^{\prime}+C_{\nu}\int_{0}^{t}\|v\|_{\dot{H}^{s}}^{2}\|v\|_{\dot{H}^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}$
since $\int_{0}^{t}\|v\|_{\dot{H}^{\frac{\gamma}{2}}}\mathrm{d}t^{\prime}$ is uniformly bounded by by constant only dependent on $\|u_{0}\|_{L^{2}}$ and $\|X\|_{C_{T}W^{\alpha}}$. By Gronwall’s inequality and $H^{s}\subset\dot{H}^{s}$ for $s\geq 0$, we prove (3.9) for all $v\in C_{T}L^{2}\cap L_{T}^{2}H^{\frac{\gamma}{2}}$. ∎
@@
Firstly, we claim that there exist $v$ satisfied equation (3.3) and energy estimate (3.7). In fact, choosing a function $\phi:\mathbb{R}\mapsto\mathbb{R}_{+}$, that is even, smooth, compactly supported, decreasing on $\mathbb{R}_{+}$, and such that $\phi(0)=1$, and then setting
@@
And $X^{\epsilon}$ solve the equation
@@
Let $v^{\epsilon}$ be the solution of equation
$\partial_{t}v^{\epsilon}-\Lambda^{\gamma}v^{\epsilon}=\frac{1}{2}\partial_{x}(v^{\epsilon})^{2}+\partial_{x}(v^{\epsilon}X^{\epsilon})+\frac{1}{2}\partial_{x}(X^{\epsilon})^{2},\quad v^{\epsilon}(0)=u_{0}^{\epsilon}.$ (3.12)
Then $v^{\epsilon}$ satisfied Proposition 3.2. Without loss of generality, let $\|X^{\epsilon}\|_{C_{T}W^{\alpha}}\leq\|X\|_{C_{T}W^{\alpha}}$ and $\|u^{\epsilon}_{0}\|_{H^{s}}\leq\|u_{0}\|_{H^{s}}$, we get that $v^{\epsilon}$ is uniformly bounded in $C_{T}L^{2}$. Then by Fatou’s lemma and Proposition 3.2, there exists $\psi:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ and a subsequence $v_{\phi(\epsilon)}$ satisfied
$\lim_{\epsilon\to 0}\psi(\epsilon)\to 0,\quad\lim_{\epsilon\to 0}v_{\phi(\epsilon)}\to v\quad in\>S^{\prime},\quad\|v\|_{C_{T}L^{2}}^{2}+\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}s\leq C(\|u_{0}\|_{L^{2}},\|X\|_{C_{T}W^{\alpha}}).$
To prove $v$ is the solution of (3.3), we need the strong convergence of $v_{\phi(\epsilon)}$ in $L^{2}$ to make sure that $\partial_{x}(v_{\phi(\epsilon)})^{2}$ can be weak-converge to $\partial_{x}(v^{2})$(its already valid for $\partial_{x}(X^{\epsilon})^{2}$ and $\partial_{x}(X^{\epsilon}v^{\epsilon})$). We use the Aubin-Lions-Simon Theorem in [Sim86]. It suffice to show following two conditions

### v2

@@ 3.2 Well-posedness
Observing that $v$ satisfies (3.6) with $f=\partial_{x}(v+\tilde{X})^{2}$. By Lemma 3.3 and integrating by parts, we have
$\|v(t)\|_{L^{2}}^{2}+\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq\|u_{0}\|_{L^{2}}^{2}+2\int_{0}^{t}\|v(t^{\prime})\|_{L^{2}}^{2}\mathrm{d}t^{\prime}+2\int_{0}^{t}\langle v(t^{\prime}),2\partial_{x}(v(t^{\prime})\tilde{X}(t^{\prime}))+\partial_{x}(\tilde{X}^{2}(t^{\prime}))\rangle_{0}\mathrm{d}t^{\prime}.$
By Hölder inequality, Young’s inequality, Lemma 2.1 and Lemma 2.7, there exist $\alpha\in(\frac{1}{4},\frac{1}{2})$ such that for $\frac{\gamma}{2}-\frac{1}{2}>\alpha>1-\frac{\gamma}{2}$, we have estimate
$\langle v,2\partial_{x}(v\tilde{X})+\partial_{x}(\tilde{X}^{2})\rangle_{0}\leq$ $2\|v\|_{H^{\frac{\gamma}{2}}}\|v\tilde{X}\|_{H^{1-\frac{\gamma}{2}}}+\|v\|_{H^{\frac{\gamma}{2}}}\|\tilde{X}^{2}\|_{H^{1-\frac{\gamma}{2}}}$
$\leq$ $2\|v\|_{H^{\frac{\gamma}{2}}}^{1+\frac{1}{\gamma}}\|v\|_{L^{2}}^{1-\frac{1}{\gamma}}\|\tilde{X}\|_{W^{\alpha}}+\|v\|_{H^{\frac{\gamma}{2}}}\|\tilde{X}\|_{W^{\alpha}}^{2}$
$\leq$ $\frac{\nu}{2}\|v\|_{H^{\frac{\gamma}{2}}}^{2}+C_{\nu}(\|v\|_{L^{2}}^{2}\|\tilde{X}\|^{\frac{2\gamma}{\gamma-1}}_{W^{\alpha}}+\|\tilde{X}\|_{W^{\alpha}}^{4}).$
For $0\leq t\leq T$, we have
$\|v(t)\|_{L^{2}}^{2}+\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq$ $\|u_{0}\|_{L^{2}}^{2}+\nu\int_{0}^{t}\|v(t^{\prime})\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}+C_{\nu}\int_{0}^{t}(1+\|\tilde{X}(t^{\prime})\|_{W^{\alpha}}^{\frac{\gamma}{\gamma-1}})\|v(t^{\prime})\|_{L^{2}}^{2}\mathrm{d}t^{\prime}$ (3.8)
$+C_{\nu}\int_{0}^{t}\|\tilde{X}(t^{\prime})\|_{W^{\alpha}}^{4}\mathrm{d}t^{\prime}.$
Observing that $\|\tilde{X}\|_{C_{t}W^{\alpha}}\leq\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}$, for $\nu<1$, by Gronwall’s inequality, the proposition is proved. ∎
Furthermore, the following lemma shows the estimate of $v$ can be improved to higher regularity.
@@
Let $\gamma\in(\frac{3}{2},2]$, $\alpha\in(s+1-\frac{\gamma}{2},\frac{\gamma}{2}-\frac{1}{2})$ , $s<\gamma-\frac{3}{2}$, $\delta$ satisfy Theorem 1.1 and $0\leq T\leq\bar{T}<\infty$, let $v$ be the smooth solution of (3.3) on $[0,T]$ for $u_{0}\in H^{\frac{1}{2}+\delta}$ with $\tilde{X}\in C_{\bar{T}}W^{\alpha}$. For $t\in[0,T]$, we have the following energy estimate
$\|\Lambda^{s}v\|_{L^{2}}+2(1-\nu)\int_{0}^{t}\|\Lambda^{s}v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}\leq$ $\|u_{0}\|^{2}_{H^{s}}e^{C_{\nu}\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}}+C_{\nu}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{2}\int_{0}^{t}e^{C_{\nu}\int_{t^{\prime}}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime\prime}}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}$ (3.9)
$+C_{\nu}\|\tilde{X}\|_{C_{\bar{T}}W^{\alpha}}^{4}\int_{0}^{t}e^{C_{\nu}\int_{t^{\prime}}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime\prime}}\mathrm{d}t^{\prime},$
@@
Recalling $\Lambda^{s}v$ satisfies the following equation
$\partial_{t}(\Lambda^{s}v)-\Lambda^{\gamma}(\Lambda^{s}v)=\Lambda^{s}\partial_{x}(v^{2})+2\Lambda^{s}\partial_{x}(v\tilde{X})+\Lambda^{s}\partial_{x}(\tilde{X}^{2}),$ (3.10)
multiplying $\Lambda^{s}v$ both sides and integrating on $\mathbb{T}$, we have
$\frac{1}{2}\partial_{t}\|\Lambda^{s}v\|_{L^{2}}^{2}+\|\Lambda^{s}v\|_{\dot{H}^{\frac{\gamma}{2}}}^{2}=\langle\Lambda^{s}\partial_{x}(v^{2}),\Lambda^{s}v\rangle+2\langle\Lambda^{s}\partial_{x}(v\tilde{X}),\Lambda^{s}v\rangle+\langle\Lambda^{s}\partial_{x}(\tilde{X}^{2}),\Lambda^{s}v\rangle.$ (3.11)
For the first term in the right side of (3.11), by Parsavel formula and Hölder inequality, we have
$\langle\Lambda^{s}\partial_{x}(v^{2}),\Lambda^{s}v\rangle\leq$ $\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}\|D^{{s+1-\frac{\gamma}{2}}}(v^{2})\|_{L^{2}}$
$\leq$ $\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}\|D^{{s+1-\frac{\gamma}{2}}}v^{2}-2vD^{s+1-\frac{\gamma}{2}}v\|_{L^{2}}+\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}\|2vD^{s+1-\frac{\gamma}{2}}v\|_{L^{2}}.$
By the estimate in [KPV93, GO14] and the interpolation, for $0<s+1-\frac{\gamma}{2}<1$ and $\alpha=\frac{s}{2}+\frac{1}{2}-\frac{\gamma}{4}$, we have
$\|D^{{s+1-\frac{\gamma}{2}}}v-2vD^{s+1-\frac{\gamma}{2}}v\|_{L^{2}}\lesssim$ $\|D^{\beta}v\|_{L^{4}}^{2}$
@@
for $\gamma\geq\frac{3}{2}$. By Hölder’s inequality and Sobolev embedding in [BCD11], we have
$\|2vD^{s+1-\frac{\gamma}{2}}v\|_{L^{2}}\lesssim$ $\|v\|_{L^{\frac{2}{1-2s}}}\|D^{s+1-\frac{\gamma}{2}}v\|_{L^{\frac{2}{2s}}}$
$\lesssim$ $\|v\|_{\dot{H}^{s}}\|D^{s+1-\frac{\gamma}{2}}v\|_{\dot{H}^{\frac{1}{2}-s}}$
$\lesssim$ $\|v\|_{\dot{H}^{s}}\|v\|_{\dot{H}^{\frac{3}{2}-\frac{\gamma}{2}}}$
$\lesssim$ $\|v\|_{\dot{H}^{s}}\|v\|_{H^{\frac{\gamma}{2}}},$
@@
$\langle\Lambda^{s}\partial_{x}(v^{2}),\Lambda^{s}v\rangle\leq\frac{\nu}{3}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}^{2}+C_{\nu}\|v\|_{\dot{H}^{s}}^{2}\|v\|_{\dot{H}^{\frac{\gamma}{2}}}^{2}.$
For the second term and third term of right hand side of (3.11), by Parsavel formula Hölder’s inequality and Lemma 2.1, for $s+1-\frac{\gamma}{2}\leq\alpha<\frac{\gamma}{2}-\frac{1}{2}$, we have
$2\langle\Lambda^{s}\partial_{x}(vX),\Lambda^{s}v\rangle\leq$ $\|vX\|_{\dot{H}^{s+1-\frac{\gamma}{2}}}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}$
$\leq$ $\|X\|_{W^{\alpha}}\|v\|_{H^{s+\frac{3}{2}-\frac{\gamma}{2}}}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}$
$\leq$ $\|X\|_{W^{\alpha}}\|v\|_{H^{\frac{\gamma}{2}}}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}$
$\leq$ $C_{\nu}\|X\|_{W^{\alpha}}^{2}\|v\|^{2}_{H^{\frac{\gamma}{2}}}+\frac{\nu}{3}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}^{2},$
@@
$\langle\Lambda^{s}\partial_{x}(X^{2}),\Lambda^{s}v\rangle\leq$ $\|X^{2}\|_{\dot{H}^{s+1-\frac{\gamma}{2}}}\|v\|_{\dot{H}^{s+\frac{\gamma}{2}}}$
@@
Combining the above estimates, and integral (3.11) on $[0,t]$, we have
$\frac{1}{2}\|\Lambda^{s}v(t)\|_{L^{2}}^{2}+(1-\nu)\int_{0}^{t}\|\Lambda^{s}v(t^{\prime})\|_{\dot{H}^{\frac{\gamma}{2}}}\mathrm{d}t^{\prime}\leq$ $\frac{1}{2}\|u_{0}\|_{H^{s}}^{2}+C_{\nu}\int_{0}^{t}\|X(t^{\prime})\|_{W^{\alpha}}^{2}\|v(t^{\prime})\|_{\dot{H}^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}+C_{\nu}\int_{0}^{t}\|X(t^{\prime})\|_{W^{\alpha}}^{4}\mathrm{d}t^{\prime}$
$+C_{\nu}\int_{0}^{t}\|v(t^{\prime})\|_{\dot{H}^{s}}^{2}\|v(t^{\prime})\|_{\dot{H}^{\frac{\gamma}{2}}}^{2}\mathrm{d}t^{\prime}.$
Since $\int_{0}^{t}\|v\|_{\dot{H}^{\frac{\gamma}{2}}}\mathrm{d}t^{\prime}$ is uniformly bounded by constant depending only on $\|u_{0}\|_{L^{2}}$ and $\|X\|_{C_{\bar{T}}W^{\alpha}}$. By Gronwall’s inequality and $H^{s}\subset\dot{H}^{s}$ for $s\geq 0$, we prove (3.9) for all $v\in C_{T}L^{2}\cap L_{T}^{2}H^{\frac{\gamma}{2}}$. ∎
@@
Firstly, we claim that there exist $v$ satisfying equation (3.3) and energy estimate (3.7). In fact, choosing a function $\phi:\mathbb{R}\mapsto\mathbb{R}_{+}$, that is even, smooth, compactly supported, decreasing on $\mathbb{R}_{+}$, and such that $\phi(0)=1$, and then setting
@@
And $X^{\epsilon}$ solves the equation
@@
Without loss of generality, we assume that $\|X^{\epsilon}\|_{C_{T}W^{\alpha}}\leq\|X\|_{C_{T}W^{\alpha}}$ and $\|u^{\epsilon}_{0}\|_{H^{s}}\leq\|u_{0}\|_{H^{s}}$. Let $v^{\epsilon}$ be the solution of equation
$\partial_{t}v^{\epsilon}-\Lambda^{\gamma}v^{\epsilon}=\partial_{x}(v^{\epsilon})^{2}+2\partial_{x}(v^{\epsilon}X^{\epsilon})+\partial_{x}(X^{\epsilon})^{2},\quad v^{\epsilon}(0)=u_{0}^{\epsilon}.$ (3.12)
and $T^{*}$ be the time given by Theorem 1.1. Then $v^{\epsilon}$ satisfies the condition of Proposition 3.2 on $0,T^{*}]$. Noting that $X\in C_{T}W^{\alpha}$ for $s+1-\frac{\gamma}{2}<\alpha<\frac{\gamma}{2}-\frac{1}{2}$ and $s<\gamma-\frac{3}{2}$ and any $T>0$. We get that $v^{\epsilon}$ is uniformly bounded in $C_{T^{*}}L^{2}$. Then by Banach–Alaoglu Theorem, Fatou’s lemma([BCD11]) and Proposition 3.2, there exists $\psi:\mathbb{R}_{+}\mapsto\mathbb{R}_{+}$ and a subsequence $v_{\phi(\epsilon)}$ satisfying
$\lim_{\epsilon\to 0}\psi(\epsilon)\to 0,\quad\lim_{\epsilon\to 0}v_{\phi(\epsilon)}\to v\quad in\>L^{2},\quad\|v\|_{C_{T}L^{2}}^{2}+\int_{0}^{t}\|v\|_{H^{\frac{\gamma}{2}}}^{2}\mathrm{d}s\leq C(\|u_{0}\|_{L^{2}},\|X\|_{C_{T}W^{\alpha}}).$
To prove $v$ is the solution of (3.3), we need the strong convergence of $v_{\phi(\epsilon)}$ in $L^{2}$ to make sure that $\partial_{x}(v_{\phi(\epsilon)})^{2}$ can be weak-converge to $\partial_{x}(v^{2})$(it’s already valid for $\partial_{x}(X^{\epsilon})^{2}$ and $\partial_{x}(X^{\epsilon}v^{\epsilon})$). We use the Aubin-Lions-Simon Theorem in [Sim86]. It is suffice to show the following two conditions